OpenStudy (anonymous):

. Let z = (sq rt 3) – i and w = -2 – 2 i . Calculate z^2. Express your answer in rectangular (a + bi) form.

OpenStudy (anonymous):

z^2 = 3-1-2sqrt3 x i = 2-2sqrt3 i

OpenStudy (anonymous):

@sachinrajsharma caculate w^4 express answer in polar (r(costheta+isintheta)

OpenStudy (anonymous):

w^2 =-( 4-4-8i) = 8i w^4 = (8i)^2 = -64 = -64+0i r = sqrt(64^2-0i^2) = 64 arctan(y/x) =arctan(0/64) = 0 value of theta = 0 => polar form is = 64(cos 0 +isin0)